OpenStudy (anonymous):

Find the product of (x + 5)(x - 5). x2 - 10x + 25 x2 + 10x + 25 x2 - 25 x2 + 25 @Directrix @amistre64 @ganeshie8

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

If you FOIL \[x \times x=x^2\]\[x \times -5=-5x\]\[x \times 5=5x\]\[-5 \times 5=-25\] So you'll get \[x^2-5x+5x-25\]

OpenStudy (anonymous):

so its A?

OpenStudy (anonymous):

@wrstlr3232

OpenStudy (anonymous):

Not A. because you have the -5x and 5x. Since they both have an x to the 1 power, you can combine those. So -5x+5x=?

OpenStudy (anonymous):

d or c?

OpenStudy (anonymous):

so its C @wrstlr3232

OpenStudy (anonymous):

C, correct.